# Object Detection for Remote Sensing Community

### v1

@@ Abstract
Object detection, particularly open-vocabulary object detection, plays a crucial role in Earth sciences, such as environmental monitoring, natural disaster assessment, and land-use planning. However, existing open-vocabulary detectors, primarily trained on natural-world images, struggle to generalize to remote sensing images due to a significant data domain gap. Thus, this paper aims to advance the development of open-vocabulary object detection in remote sensing community. To achieve this, we first reformulate the task as Locate Anything on Earth (LAE) with the goal of detecting any novel concepts on Earth. We then developed the LAE-Label Engine which collects, auto-annotates, and unifies up to 10 remote sensing datasets creating the LAE-1M — the first large-scale remote sensing object detection dataset with broad category coverage. Using the LAE-1M, we further propose and train the novel LAE-DINO Model, the first open-vocabulary foundation object detector for the LAE task, featuring Dynamic Vocabulary Construction (DVC) and Visual-Guided Text Prompt Learning (VisGT) modules. DVC dynamically constructs vocabulary for each training batch, while VisGT maps visual features to semantic space, enhancing text features. We comprehensively conduct experiments on established remote sensing benchmark DIOR, DOTAv2.0, as well as our newly introduced 80-class LAE-80C benchmark. Results demonstrate the advantages of the LAE-1M dataset and the effectiveness of the LAE-DINO method. All the datasets and codes will be available at https://github.com/jaychempan/LAE-DINO.
@@ Introduction
As one of the most fundamental and important tasks in the file of computer vision, object detection (OD) (Ren et al. 2015) has been extensively studied over the years, leading to the development of numerous detectors. In particular, open-vocabulary object detection (OVD) (Zareian et al. 2021) has been receiving increasing attention. OVD relaxes the limitation of close-set object categories in the traditional OD, allowing the detection of any novel concept during the testing time. Among various OVD methods, DINO (Zhang et al. 2023) based detectors, e.g., GroundingDINO (Liu et al. 2024), have recently shown promising performance on mainstream OVD benchmarks.
However, almost all of the state-of-the-art OVD methods are trained and tested on natural-world images. When applied to Earth science-related tasks, such as environmental monitoring (Himeur et al. 2022), natural disaster assessment (Albahri et al. 2024), land-use planning (Chaturvedi et al. 2021), these methods struggle to generalize due to the huge data domain gap. Unlike natural-world imagery, Earth science relies on remote sensing imagery, which exhibit much higher resolutions, distinct image styles (Pan et al. 2023b), and different semantic class concepts. This makes the direct transfer of current OVD models nontrivial. Therefore, in this paper, we are motivated to advance open-vocabulary object detection for remote sensing community.
@@
To achieve this goal, we first reformulate the task of OVD for remote sening filed as Locate Anything on Earth (LAE). As illustrated in Figure 1, our aim is to enable LAE models could detect any novel concept on Earth. Our efforts are mainly made from two key aspects: first, a LAE-Label Engine is developed to construct the large-scale remote sensing training data; second, a novel LAE-DINO Model is proposed and trained to work as the first foundation models for the newly proposed LAE task.
@@
Technically, the LAE-DINO, a DINO-based OVD method, is proposed and trained on the LAE-1M dataset. The novel modules of LAE-DINO are designed to address two questions: 1) How to fit the OVD model in the training data that has around 1600 vocabularies? 2) How can the relationship between image and text be better utilized to achieve more effective vocabulary-conditioned object detection? As the answer of the first question, the Dynamic Vocabulary Construction (DVC) which dynamically selects the positive and negative vocabularies for each training batch is proposed. While the Visual-Guided Text Prompt Learning (VisGT) is presented to address the second issue. Based one the observation that different objects within a single image collectively define the scene, VisGT introduces the concept of “scene features” by averaging all object features. Through taking the scene features as a bridge, VisGT aligns visual features with text features, thereby enhancing the interaction between these two modalities.
Extensive experiments are conducted on both open-set and close-set scenarios. Different models are compared taking different data as training data. Results reveal: 1) our proposed LAE-1M dataset significantly improves model performance, especially in open-set scenarios; and 2) our LAE-DINO model achieves state-of-the-art performance.
@@
Generic Object Detection for Remote Sensing
@@
Object detection (OD) is one of the classical vision tasks in computer vision, which is to obtain the locations of regions of interest from a given image. Modern object detection methods can be divided into single-stage and two-stage object detection. Single-stage methods (e.g. YOLO Family (Redmon and Farhadi 2017)) perform classification and regression directly on a predefined mass of anchor boxes. The two-stage methods (e.g. Faster R-CNN (Ren et al. 2015)) fine-tune bounding boxes based on the single-stage, which is usually more accurate than the single-stage method but is slower. Some representative work, such as DINO-based detector (Zhang et al. 2023) based on Transformer, explore the trade-off between performance and computational cost for more robust detectors. In the remote sensing community, extensive research efforts (Mall et al. 2023; Tao et al. 2023; Cong et al. 2022; Reed et al. 2023; Bastani et al. 2023; Sun et al. 2022a; Guo et al. 2024) are concentrated on the extraction of fundamental imagery knowledge from large volumes of unlabeled data, utilizing advanced self-supervised or unsupervised methodologies. Some methods (e.g. CALNet (He et al. 2023)) work on involving visible (RGB) and infrared (IR) images to enhance detection performance. While these methods are broadly applicable, they exhibit limited effectiveness in enhancing detection capabilities.
Object Detection from Few-Shot Learning to Open-Vocabulary Learning
@@
Few-Shot Object Detection (FSOD) (Chen et al. 2018) emulates the human ability to recognize unseen objects with minimal guidance by identifying and localizing new objects in images based on a query template. FSOD approaches are divided into fine-tuning-based (Chen et al. 2018), which transfers knowledge (Hospedales et al. 2021) from base to novel classes, and meta-learning-based, which uses “learning to learn” to generalize across novel classes. Recent research (Fu et al. 2024) has begun exploring cross-domain FSOD, but improvements are limited, particularly as current methods depend heavily on the quality of base classes rather than novel classes (which are from the base class).
@@
Therefore, Open-Vocabulary Object Detection (OVD) adopts a more practice-oriented learning paradigm  [^1] OVD relaxes the stringent definition of novel classes from “not seen” to “not seen finely” with weakly supervised learning (Zareian et al. 2021). compared with FSOD, aiming to construct an open visual-semantic space to enhance out-of-category identification and localisation. OVR-CNN (Zareian et al. 2021) first proposed to acquire knowledge from natural language vocabularies by pre-training the backbone with image-caption data. After that, RegionCLIP (Zhong et al. 2022) and GLIP (Li et al. 2022) became unified with the image-text matching task, expanding the visual-semantic space with more powerful flooding capabilities. While the previous work mainly improves zero-shot recognition with the help of vision-language pre-training, Grounding-DINO (Liu et al. 2024) obtained a more robust grounding capability by introducing a stronger detector structure and fine-grained multimodal feature fusion. CasDet (Li et al. 2024) combines semi-supervised learning and OVD to augment aerial detection. Due to insufficient domain annotation data, these works are weaker in open-set detection, although some show promising results in closed-set detection.
@@
To facilitate the development of generalizable detectors in the remote sensing community, we first define the task of Locate Anything on Earth (LAE) and then build a data engine, LAE-Label Engine, to construct the large-scale training datasets for learning foundation LAE models.
@@
LAE draws inspiration from the Open-Vocabulary Object Detection (OVD) task but is specifically tailored for the remote sensing field. Given remote-sensing imagery as input, LAE aims to achieve robust object recognition and localization based on provided text prompts. LAE maintains a base training dataset $\mathcal{D}_{base}$ and any potential testing dataset $\mathcal{D}_{test}$. Formally, the base dataset is represented as $\mathcal{D}_{base}=\{I,\{(b,y)_{r}\}\}$, where $I$ denotes a remote sensing image, and each image comprises $r$ objects with corresponding localization annotations $b$ and category annotations $y$. Specifically, $I$ is defined as $I\in\mathbb{R}^{H\times W\times C}$, $b$ as $b\in\mathbb{R}^{4}$, and $y$ as an element of $\mathcal{V}_{base}$, where $\mathcal{V}_{base}$ is the set of vocabularies present in $\mathcal{D}_{base}$. A large $\mathcal{V}_{base}$ is generally preferable for training foundational LAE models effectively. Moreover, we define $\mathcal{V}_{\Omega}$ as the entire language vocabulary and $\mathcal{V}_{test}$ as the testing vocabulary within $\mathcal{D}_{test}$. Consistent with the fundamental settings of OVD (Zareian et al. 2021), no constraints are imposed on $\mathcal{V}_{test}$, indicating that it can be any subset of $\mathcal{V}_{\Omega}$.
@@
To tackle these two limitations, we propose the LAE-Lable data engine which makes use of both the well-labeled data and the massive unlabeled data. More specifically, as shown in Figure 2(a), for those well-labeled datasets, we first slice the huge image of different datasets and then unify the format as the same. This part results in our fine-grained LAE-FOD dataset; For the unlabeled data, as in Figure 2(b), we build a comprehensive semi-automated data construction flow based on SAM (Kirillov et al. 2023) and Large Vision-Language Model (LVLM). We begin by extracting the location information of Regions of Interest (RoIs) from remote sensing seed datasets using SAM. The detailed information of the seed datasets is listed in Table 1. Next, we obtain the categories of the zoomed-in ROI areas by taking advantage of the LVLM, i.e., InternVL (Chen et al. 2024) which has a powerful zero-shot recognition as learned on huge amounts of data with the text prompt as demonstrated in Figure 2(b). Finally, we filter out invalid and irrelevant categories using a rule-based method. In this way, our coarse-grained LAE-COD dataset is constructed, offering a rich vocabulary for open-vocabulary pre-training.
Figure 2: The pipeline of our LAE-Label Engine.
Figure 3: The pipeline for solving the LAE task: LAE-Label Engine expands vocabulary for open-vocabulary pre-training; LAE-DINO is a DINO-based open-vocabulary detector with Dynamic Vocabulary Construction (DVC) and Visual-Guided Text Prompt Learning (VisGT), which has a pre-training and fine-tuning paradigm for open-set and closed-set detection.
@@
Due to the huge success of DINO (Zhang et al. 2023), the recent DINO-based detector e.g., GroudingDINO (Liu et al. 2024) and VideoGrounding-DINO (Wasim et al. 2024), show promising detection performance on open-vocabulary detection scenarios. Thus, in this paper, we also build our method upon the DINO and form our novel LAE-DINO detector. As illustrated in Figure 3, except for the data engine part, our LAE-DINO mainly contains the Dynamic Vocabulary Construction (DVC), the Image Backbone $E_{img}$, the Text Backbone $E_{text}$, the Visual-Guided Text Prompt Learning (VisGT), the Transformer Encoder $E_{T_{E}}$, the Query Selection $M_{qs}$, the Transformer Decoder $E_{T_{D}}$, and the Detection Head $M_{det}$. Note that the $E_{img}$, $E_{text}$, $E_{T_{E}}$, $M_{qs}$, $E_{T_{D}}$, and $M_{det}$ are basic and common modules in DINO-based detectors, thus we keep them same with the former GroudingDINO. While the DVC and the VisGT are newly proposed in this paper. Typically, the DVC is proposed to tackle the large vocabulary set issue posed by our constructed training data, and the VisGT is a novel method that uses the visual information to further guide and transform the text features.
@@
Though developed in different directions and with different new modules, the DINO-based detectors basically share the same core pipeline: Given the training dataset $\mathcal{D}_{base}$, the first thing is to construct the vocabulary set $\mathcal{V}_{base}$ by simply merging all the existing vocabularies. The vocabulary set includes positive vocabularies for categories in the images and negative words for those not seen during training. For each batch training iteration, as indicated in Figure 3, the image backbone $E_{img}$ and the text backbone $E_{text}$ are used to extract the visual features $F_{I}\in\mathbb{R}^{n_{I}\times d}$, the text features $F_{T}\in\mathbb{R}^{n_{T}\times d}$ from the input image $I$ and vocabulary set $\mathcal{V}_{base}$, respectively. The $n_{I}$ and $n_{T}$ mean the number of image and text tokens, while the $d$ denotes the dimension of features. Usually, the Swin-Transformer (Liu et al. 2021) is used as the $E_{img}$ and the BERT (Devlin et al. 2018) is used as the $E_{text}$. In addition, since the $\mathcal{V}_{base}$ contains both the positive and negative vocabularies, we further denote the text features generated from $n_{T_{p}}$ positive vocabularies as $F_{T}^{P}=[\tilde{F}_{T_{1}},\tilde{F}_{T_{2}},...,\tilde{F}_{T_{p}}]\in\mathbb{R}^{n_{T_{p}}\times d}$.
After that, the Transformer encoder $E_{T_{E}}$ which takes both the image features $F_{I}$ and the text features $F_{T}$ are applied to fuse the multi-modal features. Then, the query selection $M_{qs}$ is used to initialize the region queries which consists of the learnable content queries (Miao et al. 2023) and dynamic positional queries (Liu et al. 2022). Finally, the Transformer decoder $E_{T_{D}}$ and the detection head $M_{det}$ output the both the location and category predictions $\{(\hat{b},\hat{y})_{r}\}$ for modality alignment.
Upon the predictions $\{(\hat{b},\hat{y})_{r}\}$ and the ground truth $\{(b,y)_{r}\}$, two classical losses are calculated. One is the standard Cross Entropy (CE) loss $\mathcal{L}_{cls}$ (Li et al. 2022; Liu et al. 2024) for evaluating the classification results between $\hat{y}$ and $y$, another is the Generalized Intersection over Union (GIoU) loss $\mathcal{L}_{loc}$ (Rezatofighi et al. 2019) for evaluating the locations. The detailed calculation method for $\mathcal{L}_{cls}$ and $\mathcal{L}_{loc}$ are as follows,
@@
Constructing the vocabulary is the necessary step for OVD models, the common practise in previous work (Li et al. 2022; Liu et al. 2024) simply merges all categories directly as a vocabulary which is also the most natural way. However, such a way is nontrivial when the training dataset has a large number of vocabularies. In our case, our proposed benchmark has around 1600 vocabulary which is much more than the current language could handle at once. APE (Shen et al. 2024) try to blend the individual concepts of vocabularies as independent text prompts but discarding the correlation among vocabularies. To solve this problem, the dynamic vocabulary construction (DVC) is proposed by us. Our DVC sets a dynamic vocabulary length $N_{\mathcal{DV}}$, for each training iteration, several positive and negative vocabularies will be selected to form the $N_{\mathcal{DV}}$ vocabulary set. Concretely, if the number of base vocabulary $\mathcal{V}_{base}$ is larger than $N_{\mathcal{DV}}$, i.e., $\|\mathcal{V}_{base}\|>N_{\mathcal{DV}}$, we set the vocabularies that contained in the input image as positive ones, while randomly select the resting part from the non-postive vocabulary candidates. Otherwise, we directly keep $\mathcal{V}_{base}$ as whole set which naturally contains positive vocabularies and negative ones.
@@
OVD models primarily reply on the relationship between image and text to achieve the open-vocabulary learning. Current DINO-based detectors, including our LAE-DINO, utilize this relations through textual prompt learning. However, a picture paints a thousand word which means that sparse and limited categories are hard to fully represent a image. Also inspired by MQ-Det (Xu et al. 2024), incorporating visual prompts from additional supported images with text prompts, we propose the VisGT module which aims at leveraging the visual information to further improve the semantic representation. Notably our VisGT does not utilise visual prompts like MQ-Det, but rather visual-guided text prompts to compensate for the lack of single text prompt. Specifically, as in Figure 3, VisGT is not an object-level alignment but an image-level alignment that represents the overall objects of the scene, preserving the knowledge of vocabulary to fine-grained detection across different categories.
The detailed architecture of our VisGT in shown in Figure 4. First of all, we propose the “scene features” by fusing different text features. The observation behind this is that we think the different object categories together could convert some useful scene information. For example, airplane and vehicle are two typical concepts that are strongly related to the airport scene. Thus, given the textual features $F_{T}$ with its positive textual features $F_{T_{P}}$, we define the scene feature $s$ as,
@@
where $L_{i}$ is the token length of the $i$-th category, which corresponds to the $T_{i}$-th token. By combining different instance-agnostic positive text features $F_{T}^{P}$, our scene features $s$ could be regarded as some special feature that contains both the instance-level and category-relative features. This scene feature $s$ works as the ground truth when we try to map the visual information also into the semantic space.
@@
where $FFN^{l}(\cdot)$ is the $l$-th layer of the Feed-Forward Network, and $MDSA^{l}(\cdot)$ represents the $l$-th of the MDSA module. We denote the transformed visual features as $\hat{\mathcal{S}}_{v2t}$ where the “v2tv2t” shows that our expectation of transferring the feature from visual space to textural space. Suppose that we have already learned good $\hat{\mathcal{S}}_{v2t}$, to facilitate the enhancement of visual and textural features, we combine the original text features $F_{T}$ together with the $\hat{\mathcal{S}}_{v2t}$ as the input of the Transformer encoder $E_{T_{E}}$ as,
@@
where $\tau$ is the temperature parameter and ${\phi}_{i,j}=\hat{s}_{i}^{l\mathrm{T}}s_{j}$ denotes the similarity matrix. with the $\mathcal{L}_{VisGT}$ and the classical classification loss $\mathcal{L}_{cls}$, localization loss $\mathcal{L}_{loc}$, our final loss function is as,
@@
This section first presents the experimental setup and then shows the detection results and ablation studies.

### v3

@@ Abstract
Object detection, particularly open-vocabulary object detection, plays a crucial role in Earth sciences, such as environmental monitoring, natural disaster assessment, and land-use planning. However, existing open-vocabulary detectors, primarily trained on natural-world images, struggle to generalize to remote sensing images due to a significant data domain gap. Thus, this paper aims to advance the development of open-vocabulary object detection in remote sensing community. To achieve this, we first reformulate the task as Locate Anything on Earth (LAE) with the goal of detecting any novel concepts on Earth. We then developed the LAE-Label Engine which collects, auto-annotates, and unifies up to 10 remote sensing datasets creating the LAE-1M — the first large-scale remote sensing object detection dataset with broad category coverage. Using the LAE-1M, we further propose and train the novel LAE-DINO Model, the first open-vocabulary foundation object detector for the LAE task, featuring Dynamic Vocabulary Construction (DVC) and Visual-Guided Text Prompt Learning (VisGT) modules. DVC dynamically constructs vocabulary for each training batch, while VisGT maps visual features to semantic space, enhancing text features. We comprehensively conduct experiments on established remote sensing benchmark DIOR, DOTAv2.0, as well as our newly introduced 80-class LAE-80C benchmark. Results demonstrate the advantages of the LAE-1M dataset and the effectiveness of the LAE-DINO method.
Codehttps://github.com/jaychempan/LAE-DINO
@@ Introduction
As one of the most fundamental and important tasks in the file of computer vision, object detection (OD) and localization  (Ren et al. 2015) has been extensively studied over the years, leading to the development of numerous detectors. In particular, open-vocabulary object detection (OVD) (Zareian et al. 2021) has been receiving increasing attention. OVD relaxes the limitation of close-set object categories in the traditional OD, allowing the detection of any novel concept during the testing time. Among various OVD methods, DINO (Zhang et al. 2023) based detectors, e.g., GroundingDINO (Liu et al. 2024b), have recently shown promising performance on mainstream OVD benchmarks.
However, almost all of the state-of-the-art OVD methods are trained and tested on natural-world images. When applied to Earth science-related tasks, such as environmental monitoring, natural disaster assessment, land-use planning, these methods struggle to generalize due to the huge data domain gap. Unlike natural-world imagery, Earth science relies on remote sensing imagery, which exhibit much higher resolutions, distinct image styles (Ma, Pan, and Bai 2024; Pan et al. 2023b), and different semantic class concepts. This makes the direct transfer of current OVD models nontrivial. Therefore, in this paper, we are motivated to advance open-vocabulary object detection for remote sensing community.
@@
To achieve this goal, we first reformulate the task of OVD for remote sening filed as Locate Anything on Earth (LAE)111 Following (Zhang and Wang 2024), we use “localization” to describe detection tasks in the remote sensing domain. . As illustrated in Figure 1, our aim is to enable LAE models could detect any novel concept on Earth. Our efforts are mainly made from two key aspects: first, a LAE-Label Engine is developed to construct the large-scale remote sensing training data; second, a novel LAE-DINO Model is proposed and trained to work as the first foundation models for the newly proposed LAE task.
@@
Technically, the LAE-DINO, a DINO-based OVD method, is proposed and trained on the LAE-1M dataset. The novel modules of LAE-DINO are designed to address two questions: 1) How to fit the OVD model in the training data that has around 1600 vocabularies? 2) How can the relationship between image and text be better utilized to achieve more effective vocabulary-conditioned object detection? As the answer of the first question, the Dynamic Vocabulary Construction (DVC) which dynamically selects the positive and negative vocabularies for each training batch is proposed. While the Visual-Guided Text Prompt Learning (VisGT) is presented to address the second issue. Based one the observation that different objects within a single image collectively define the scene, VisGT introduces the concept of “scene features” by averaging all object features. Through taking the scene features as a bridge, VisGT aligns visual features with text features, thereby enhancing the interaction between these two modalities. Extensive experiments are conducted on both open-set and close-set scenarios. Different models are compared taking different data as training data. Results reveal: 1) our proposed LAE-1M dataset significantly improves model performance, especially in open-set scenarios; and 2) our LAE-DINO model achieves state-of-the-art performance.
@@
Generic Object Detection for Remote Sensing.
@@
Object detection (OD) is one of the classical vision tasks in computer vision, which is to obtain the locations of regions of interest from a given image. Object detection methods can be mainly divided into single-stage and two-stage object detection. Single-stage methods (e.g. YOLO Family (Redmon and Farhadi 2017)) perform classification and regression directly on a predefined mass of anchor boxes. The two-stage methods (e.g. Faster R-CNN (Ren et al. 2015)) fine-tune bounding boxes based on the single-stage, which are usually more accurate than single-stage methods but are slower. Some representative works, such as DINO-based detector (Zhang et al. 2023) based on Transformer, explore the trade-off between performance and computational cost for more robust detectors. In the remote sensing community, extensive research efforts (Tao et al. 2023; Cong et al. 2022; Reed et al. 2023; Bastani et al. 2023; Sun et al. 2022a; Guo et al. 2024) are concentrated on the extraction of fundamental imagery knowledge from large volumes of unlabeled data, utilizing advanced self-supervised or unsupervised methodologies. Some methods (e.g. CALNet (He et al. 2023)) work on involving visible (RGB) and infrared (IR) images to enhance detection performance. While these methods are broadly applicable, they exhibit limited effectiveness in enhancing detection capabilities.
Figure 2: The pipeline of our LAE-Label Engine.
Object Detection from Few-Shot Learning to Open-Vocabulary Learning.
@@
Few-Shot Object Detection (FSOD) (Chen et al. 2018) aims to detect unseen objects using only a few labeled examples. FSOD approaches are divided into fine-tuning-based (Chen et al. 2018), which transfers knowledge (Hospedales et al. 2021) from base to novel classes, and meta-learning-based, which uses “learning to learn” to generalize across novel classes. CD-FSOD (Fu et al. 2024)) explores cross-domain FSOD (e.g., from natural to remote sensing images), yet it relies on visual images, offering limited support for new vocabulary.
@@
Therefore, Open-Vocabulary Object Detection (OVD) adopts a more practice-oriented learning paradigm (Zareian et al. 2021) compared with FSOD, aiming to construct an open visual-semantic space to enhance out-of-category identification and localisation. OVR-CNN (Zareian et al. 2021) first proposed to acquire knowledge from natural language vocabularies by pre-training the backbone with image-caption data. After that, RegionCLIP (Zhong et al. 2022) and GLIP (Li et al. 2022) became unified with the image-text matching task, expanding the visual-semantic space with more powerful flooding capabilities. While the previous work mainly improves zero-shot recognition with the help of vision-language pre-training, Grounding-DINO (Liu et al. 2024b) obtained a more robust grounding capability by introducing a stronger detector structure and fine-grained multimodal feature fusion. CasDet (Li et al. 2024b) combines semi-supervised learning and OVD to augment aerial detection. Due to insufficient domain annotation data, these works are weaker in open-set detection, although some show promising results in closed-set detection.
@@
Locate Anything on Earth (LAE) draws inspiration from the Open-Vocabulary Object Detection (OVD) task but is specifically tailored for the remote sensing field. Given remote-sensing imagery as input, LAE aims to achieve robust object recognition and localization based on provided text prompts.
LAE maintains a base training dataset $\mathcal{D}_{base}$ and any potential testing dataset $\mathcal{D}_{test}$. Formally, the base dataset is represented as $\mathcal{D}_{base}=\{I,\{(b,y)_{r}\}\}$, where $I$ denotes a remote sensing image, and each image comprises $r$ objects with corresponding localization annotations $b$ and category annotations $y$. Specifically, $I$ is defined as $I\in\mathbb{R}^{H\times W\times C}$, $b$ as $b\in\mathbb{R}^{4}$, and $y$ as an element of $\mathcal{V}_{base}$, where $\mathcal{V}_{base}$ is the set of vocabularies present in $\mathcal{D}_{base}$. A large $\mathcal{V}_{base}$ is generally preferable for training foundational LAE models effectively. Moreover, we define $\mathcal{V}_{\Omega}$ as the entire language vocabulary and $\mathcal{V}_{test}$ as the testing vocabulary within $\mathcal{D}_{test}$. Consistent with the fundamental settings of OVD (Zareian et al. 2021), no constraints are imposed on $\mathcal{V}_{test}$, indicating that it can be any subset of $\mathcal{V}_{\Omega}$.
@@
To tackle these two limitations, we propose the LAE-Lable data engine which makes use of both the well-labeled data and the massive unlabeled data. More specifically, as shown in Figure 2(a), for those well-labeled datasets, we first slice the huge image of different datasets and then unify the format as the same. This part results in our fine-grained LAE-FOD dataset; For the unlabeled data, as in Figure 2(b), we build a comprehensive semi-automated data construction flow based on SAM (Kirillov et al. 2023) and Large Vision-Language Model (LVLM).
We begin by extracting the location information of Regions of Interest from remote sensing seed datasets using SAM. The detailed information of the seed datasets is listed in Table 1. Next, we obtain the categories of the zoomed-in ROI areas by taking advantage of the LVLM, i.e., InternVL (Chen et al. 2024) which has a powerful zero-shot recognition as learned on huge amounts of data with the text prompt as demonstrated in Figure 2(b). Finally, we filter out invalid and irrelevant categories using a rule-based method. In this way, our coarse-grained LAE-COD dataset is constructed, offering a rich vocabulary for open-vocabulary pre-training.
Figure 3: The pipeline for LAE-DINO.
@@
Due to the huge success of DINO (Zhang et al. 2023), the recent DINO-based detector e.g., GroudingDINO (Liu et al. 2024b) and VideoGrounding-DINO (Wasim et al. 2024), show promising detection performance on open-vocabulary detection scenarios. Thus, in this paper, we also build our method upon the DINO and form our novel LAE-DINO detector. As illustrated in Figure 3, except for the data engine part, our LAE-DINO mainly contains the Dynamic Vocabulary Construction (DVC), the Image Backbone $E_{img}$, the Text Backbone $E_{text}$, the Visual-Guided Text Prompt Learning (VisGT), the Transformer Encoder $E_{T_{E}}$, the Query Selection $M_{qs}$, the Transformer Decoder $E_{T_{D}}$, and the Detection Head $M_{det}$. Note that the $E_{img}$, $E_{text}$, $E_{T_{E}}$, $M_{qs}$, $E_{T_{D}}$, and $M_{det}$ are basic and common modules in DINO-based detectors, thus we keep them same with the former GroudingDINO. While the DVC and the VisGT are newly proposed in this paper. Typically, the DVC is proposed to tackle the large vocabulary set issue posed by our constructed training data, and the VisGT is a novel method that uses the visual information to further guide and transform the text features.
@@
Though developed in different directions and with different new modules, the DINO-based detectors basically share the same core pipeline: Given the training dataset $\mathcal{D}_{base}$, the first thing is to construct the vocabulary set $\mathcal{V}_{base}$ by simply merging all the existing vocabularies. The vocabulary set includes positive vocabularies for categories in the images and negative words for those not seen during training.
For each batch training iteration, as indicated in Figure 3, the image backbone $E_{img}$ and the text backbone $E_{text}$ are used to extract the visual features $F_{I}\in\mathbb{R}^{n_{I}\times d}$, the text features $F_{T}\in\mathbb{R}^{n_{T}\times d}$ from the input image $I$ and vocabulary set $\mathcal{V}_{base}$, respectively. The $n_{I}$ and $n_{T}$ mean the number of image and text tokens, while the $d$ denotes the dimension of features. Usually, the Swin-Transformer (Liu et al. 2021) is used as the $E_{img}$ and the BERT (Devlin et al. 2018) is used as the $E_{text}$. In addition, since the $\mathcal{V}_{base}$ contains both the positive and negative vocabularies, we further denote the text features generated from $n_{T_{p}}$ positive vocabularies as $F_{T}^{P}=[\tilde{F}_{T_{1}},\tilde{F}_{T_{2}},...,\tilde{F}_{T_{p}}]\in\mathbb{R}^{n_{T_{p}}\times d}$.
After that, the Transformer encoder $E_{T_{E}}$ which takes both the image features $F_{I}$ and the text features $F_{T}$ are applied to fuse the multi-modal features. Then, the query selection $M_{qs}$ is used to initialize the region queries which consists of the learnable content queries and dynamic positional queries. Finally, the Transformer decoder $E_{T_{D}}$ and the detection head $M_{det}$ output the both the location and category predictions $\{(\hat{b},\hat{y})_{r}\}$ for modality alignment.
Upon the predictions $\{(\hat{b},\hat{y})_{r}\}$ and the ground truth $\{(b,y)_{r}\}$, two classical losses are calculated. One is the standard Cross Entropy (CE) loss $\mathcal{L}_{cls}$ (Li et al. 2022; Liu et al. 2024b) for evaluating the classification results between $\hat{y}$ and $y$, another is the Generalized Intersection over Union (GIoU) loss $\mathcal{L}_{loc}$ (Rezatofighi et al. 2019) for evaluating the locations. The detailed calculation method for $\mathcal{L}_{cls}$ and $\mathcal{L}_{loc}$ are as follows,
@@
Current OVD detectors (Li et al. 2022; Liu et al. 2024b) like Grounding DINO rely on fixed-token-length text encoders (e.g., BERT (Devlin et al. 2018) or CLIP (Radford et al. 2021)) that concatenate all categories into a “extremely long text”, which is nontrivial for datasets with numerous categories. For example, while BERT allows a maximum of 256 tokens, our dataset includes about 1600 categories, exceeding this limit. This motivated us to develop the dynamic vocabulary construction (DVC), which reduces the number of input categories.
To tackle with such a “extremely long text”, APE (Shen et al. 2024) tries to blend the individual concepts of vocabularies as independent text prompts but discarding the correlation among vocabularies. Our DVC sets a dynamic vocabulary length $N_{\mathcal{DV}}$, for each training iteration, several positive and negative vocabularies will be selected to form the $N_{\mathcal{DV}}$ vocabulary set. Concretely, if the number of base vocabulary $\mathcal{V}_{base}$ is larger than $N_{\mathcal{DV}}$, i.e., $\|\mathcal{V}_{base}\|>N_{\mathcal{DV}}$. In each training batch, DVC ensures the input category length is fixed at $N_{\mathcal{DV}}$. All current batch categories are considered positive categories (for example $N_{pos}$) are included in the input. The remaining ($N_{\mathcal{DV}}-N_{pos}$) part is filled by randomly sampling negative categories from the rest of the whole category set $\mathcal{V}_{base}$. DVC can effectively reduce the number of iterations for text encoder inference.
@@
OVD models primarily reply on the relationship between image and text to achieve the open-vocabulary learning. Current DINO-based detectors, including our LAE-DINO, utilize this relations through textual prompt learning.
However, a picture paints a thousand word which means that sparse and limited categories are hard to fully represent a image. Also inspired by MQ-Det (Xu et al. 2024), incorporating visual prompts from additional supported images with text prompts, we propose the VisGT module which aims at leveraging the visual information to further improve the semantic representation. Notably our VisGT does not utilise visual prompts like MQ-Det, but rather visual-guided text prompts to compensate for the lack of single text prompt.
Specifically, as in Figure 3, VisGT is not an object-level alignment but an image-level alignment that represents the overall objects of the scene, preserving the knowledge of vocabulary to fine-grained detection across different categories.
The detailed architecture of our VisGT in shown in Figure 4. First of all, we propose the “scene features” by fusing different text features. The observation behind this is that the different object categories together could convert some useful scene information. For example, airplane and vehicle are two typical concepts that are strongly related to the airport scene. Thus, given the textual features $F_{T}$ with its positive textual features $F_{T_{P}}$, we define the scene feature $s$ as,
@@
where $L_{i}$ is the token length of the $i$-th category, which corresponds to the $T_{i}$-th token.
By combining different instance-agnostic positive text features $F_{T}^{P}$, our scene features $s$ could be regarded as some special feature that contains both the instance-level and category-relative features. This scene feature $s$ works as the ground truth when we try to map the visual information also into the semantic space.
@@
where $FFN^{l}(\cdot)$ is the $l$-th layer of the Feed-Forward Network, and $MDSA^{l}(\cdot)$ means the $l$-th of the MDSA module.
We denote the transformed visual features as $\hat{\mathcal{S}}_{v2t}$ where the “$v2t$” shows that our expectation of transferring the feature from visual space to textural space.
Suppose that we have already learned good $\hat{\mathcal{S}}_{v2t}$, to facilitate the enhancement of visual and textural features, we combine the original text features $F_{T}$ together with the $\hat{\mathcal{S}}_{v2t}$ as the input of the Transformer encoder $E_{T_{E}}$ as,
@@
where $\tau$ is the temperature parameter and ${\phi}_{i,j}=\hat{s}_{i}^{l\mathrm{T}}s_{j}$ denotes the similarity matrix.
with the $\mathcal{L}_{VisGT}$ and the classical classification loss $\mathcal{L}_{cls}$, localization loss $\mathcal{L}_{loc}$, our final loss function is as,